OpenStudy (anonymous):

Derivatives: Can you tell me what I did wrong? Thanks. My questions are in the comments. *Using logarithmic differentiation

OpenStudy (anonymous):

Find the derivative of: \[\frac{ x^5cosx }{ e^(-x^2+2x) }\]

OpenStudy (anonymous):

I keep getting the answer \[(\frac{ 5 }{ x }-\frac{ sinx }{ cosx }-2+x^2)(\frac{ x^5cosx }{ e^(-x^2+2x) })\]

OpenStudy (anonymous):

But this is my teacher's answer \[(\frac{ 5 }{ x }-\frac{ sinx }{ cosx }+2x-2)(\frac{ x^5cosx }{ e^(-x^2+2x)})\]

OpenStudy (anonymous):

your teachers answer doesn't make sense quite either

OpenStudy (anonymous):

(-x^2+2x) is an exponent

OpenStudy (anonymous):

if that is the original the first one you posted you have a quotient rule along with a product rule on the top and a composition rule on the bottom

OpenStudy (anonymous):